A cylindrical container with a cross-sectional area of 70.2cm2 holds a fluid of density 796kg/m3. At the bottom of the container, their pressure is 117kpa. Assume pat=101kpa. what is the depth of the fluid? Find the pressure at the bottom Of the container after an additional 2.35×10-3m3.


1
Expert's answer
2021-11-15T17:54:32-0500

"p=p_0+\rho gh\to h=(p-p_0)/(\rho g)="


"=(117000-101000)/(796\cdot9.8)=2.051\ (m)" . Answer


"h_1=V/A=0.00235/0.00702=0.335\ (m)"


"p_1=p_0+\rho gH=p_0+\rho g (h+h_1)="


"=101000+796\cdot9.8\cdot(2.051+0.335)=119613\ (Pa)" . Answer